Two rainstorms occurred in one week in a certain area. In the first storm 30mL of rain fell per hour, and in the second storm 20mL of rain fell per hour. Rain fell that week for a total of 55 hours for a total rainfall of 1250mL . How long was each storm?




Answer by checkley77(12569)

30x+20(55-x)=125
30x+110-20x=125
10x=125-110
10x=15
x=15 hours for the 30 ml rain fall.
55-15=30 hours for the 20 ml rain fall.
Proof:
30*15+.20*35=125
45+70=125
125=125